OpenStudy (anonymous):

\(\LARGE \int_{-\infty}^{0} x 5^{-x^2} dx\) @Mimi_x3 please

OpenStudy (anonymous):

use integral by parts formula

OpenStudy (anonymous):

isn't it special in improper integrals?

OpenStudy (zarkon):

u-sub

OpenStudy (blockcolder):

First, rewrite the improper integral as a limit of an integral, then use u-sub.

OpenStudy (anonymous):

what do you mean limit of an integral @blockcolder ?

OpenStudy (blockcolder):

I mean like this: \[\Large\lim_{t \rightarrow -\infty} \int_t^0 x5^{-x^2}\ dx\]

OpenStudy (anonymous):

hmmm ohkayyyy...? then what?

OpenStudy (blockcolder):

Then use u-sub like the others said: Let u=-x^2, du=-2xdx \[\Large\lim_{t \rightarrow -\infty}\int_t^0 x5^{-x^2}\ dx=-\frac{1}{2}\lim_{t \rightarrow -\infty} \int_{-t^2}^0 5^u\ du\]

OpenStudy (lgbasallote):

wait...why is the lower limit t^2 now?

OpenStudy (lgbasallote):

hehe sorry just learning with the user here :p

OpenStudy (blockcolder):

I made a u-sub, so I plugged in the limits on x into my equation for u.

OpenStudy (anonymous):

i dont get how it becam x^2 eeither @_@ please explain further

OpenStudy (anonymous):

uhmm is it because t = x so when u became -x^2 it became -t62??

OpenStudy (anonymous):

t^2*

OpenStudy (blockcolder):

Yep. Continuing the integral: \[\Large-\frac{1}{2}\lim_{t \rightarrow -\infty} \int_{-t^2}^0 5^u\ du= -\frac{1}{2}\lim_{t \rightarrow -\infty} \left .\left ( \frac{5^u}{\ln5}\right ) \right |_{-t^2}^0\\ \Large=-\frac{1}{2 \ln5} \lim_{t \rightarrow -\infty}(1-5^{-t^2})=-\frac{1}{2 \ln5}\] because \( \Large\displaystyle \lim_{t \rightarrow \infty} 5^{-t^2}=0\).

OpenStudy (anonymous):

\(5^\infty\) is zero?? @blockcolder

OpenStudy (blockcolder):

Oops. Typo. That should be -infinity. \(\displaystyle \Large\lim_{t \rightarrow -\infty}5^{-t^2}=0\)

OpenStudy (anonymous):

so \(5^{-\infty}\) is zero?

OpenStudy (blockcolder):

Yep. :)

OpenStudy (anonymous):

is that a property?

OpenStudy (blockcolder):

Kinda. For a>1, \(\lim_{x \rightarrow -\infty}a^x=0\).

OpenStudy (anonymous):

i see..good work :D gonna practice more though